### v1

# Constraining scalar-tensor theories from higher harmonics with GW230529

Abstract
The Advanced LIGO and Virgo collaborations recently detected a gravitational wave event, GW230529_181500, during the fourth observing run, which is most plausibly attributed to the merger of a neutron star and a black hole. This observation provides an opportunity to test a class of gravitational theories that deviate from general relativity. In such theories, additional terms contribute to the gravitational wave signal only in cases of asymmetric binaries. This study focuses on two scalar-tensor models within this class of theories: the Screened Modified Gravity and Brans-Dicke theory. These models have potential applications in areas such as dark matter, dark energy, cosmic inflation, and primordial nucleosynthesis. With the GW230529_181500 and Bayesian Markov-chain Monte Carlo analyses, we derive a 90% credible lower bound as $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25.12$ by using dominant-mode correction. Asymmetric binary systems usually have a significant mass ratio, in such cases, higher harmonic modes cannot be neglected. Our work considers higher harmonic corrections from scalar-tensor theories and provides a tighter constraint of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$, with a 13.3% and 30.1% improvement respectively. Combining GW230529_181500, GW200115 and GW190814 and including higher modes, the constraint is improved to $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$. This is currently the strongest constraint from GWs, contingent upon GW190814 being an NSBH event.
1 Introduction
The Advanced LIGO [1] and Virgo [2] collaborations recently concluded the first phase of the fourth observing run (O4a) at 16:00 UTC on January 16, 2024. Shortly after the start of O4a, aLIGO recorded a gravitational wave (GW) event, GW230529_181500 (abbreviated as GW230529), which is most plausibly attributed to the merger of a neutron star and a black hole [3].
In comparison to other methodologies such as laboratory and Solar System experiments, or binary pulsar and cosmological observations, GW events offer a powerful means of testing general relativity (GR) under strong-field or highly dynamical conditions [4, 5, 6, 7, 8, 9]. Such scenarios are particularly useful for distinguishing GR from modified gravity theories. As the number of detected GW events grows, they become increasingly valuable tools for placing constraints on modified gravity frameworks, such as Einstein-dilaton-Gauss-Bonnet gravity [10, 11, 12, 13, 14], extra dimension [10], time-varying gravitational constant G(t) [15] and scalar-tensor theories [10, 16, 17].
Scalar-tensor theories [8, 18, 19], one of the most natural extensions of GR, introduce additional scalar degrees of freedom that couple nonminimally to the metric tensor in the gravitational sector. Such scalar fields can arise through compactification mechanisms from higher-dimensional theories like string theory [20], Kaluza-Klein-like models [21], or braneworld scenarios [22, 23]. Scalar-tensor frameworks offer a comprehensive platform to explore the phenomenology of fundamental theories and have been applied to study phenomena such as the Universe’s accelerated expansion [24, 25, 26], inflation [27, 28, 29], large-scale structure formation [30], and primordial nucleosynthesis [31, 32, 33, 34]. Here, we focus on two different models of scalar-tensor theories, i.e., the screened modified gravity (SMG) and the Brans-Dicke (BD) theory. Screened modified gravity theories, scalar degrees of freedom are introduced that can serve as dynamical dark energy components while evading stringent local gravity constraints through screening mechanisms [27]. These mechanisms, such as the Chameleon [35, 36], Vainshtein [37, 38], and Symmetron [39] mechanisms, enable the suppression of deviations from GR in high-density or small-scale environments, thereby ensuring consistency with precise laboratory and Solar System tests. On cosmological scales, however, the scalar field can remain unscreened and dynamically drive the late-time acceleration of the Universe, offering a unified framework for modified gravity and dark energy phenomenology. Extensive studies have explored the theoretical consistency, astrophysical consequences, and cosmological signatures of Screened modified gravity theories [40, 27, 41]. For Brans-Dicke theory [42], originally developed by Jordan, Fierz, Brans, and Dicke, which incorporates Mach’s principle by promoting the gravitational constant to a dynamical scalar field. The interaction strength is governed by a dimensionless coupling parameter $\omega_{\rm BD}$, with general relativity recovered in the limit $\omega_{\rm BD}\to\infty$. Extensive observational efforts have placed bounds on $\omega_{\rm BD}$ using various systems. For instance, binary pulsar timing has constrained $\omega_{\rm BD}>25000$ from the orbital decay of PSR J1738+0333 [43], while the most stringent limit to date, $\omega_{\rm BD}>40000$, arises from the Shapiro time delay measured by the Cassini mission in the Solar System [44, 7].
Meanwhile, there are also many other applications for GW in constraints. In scalar-tensor theory, the waveform are influenced by neutron star–black hole (NSBH) binaries because the deviations from GR in the waveform are related to the ‘sensitivities’ of the binary components [43]. Therefore, current limits on scalar-tensor theory are rely on NSBH systems, such as using space-based GW detectors in future simulations [45, 46, 47, 48, 49]. In analyses based on observed gravitational wave events, Rui Niu et al. analyzed the GW200115 event and derived the bounds $\frac{\varphi_{\mathrm{VEV}}}{M_{\mathrm{Pl}}}<1.8\times 10^{-2}$ and $\omega_{\mathrm{BD}}>40$ [17] using Bayesian inference implemented with the open-source package Bilby [50]. Tan et al. conducted a joint analysis of GW200115 and GW190814 and found that $\frac{\varphi_{\mathrm{VEV}}}{M_{\mathrm{Pl}}}<8.32\times 10^{-3}$ and $\omega_{\mathrm{BD}}>110.55$ [51]. Similarly, using the breathing scalar polarization, Takeda et al. reported a constraint of $\omega_{\rm BD}>81$ [52].
The observational sources currently available for constraining scalar-tensor theory are largely limited to asymmetric binary systems, including NSBH, white dwarf–neutron star, and white dwarf–black hole binaries. GW230529 provides a valuable new data point for this purpose. Our work incorporates higher harmonic modes into the waveform, as such effects cannot be neglected in asymmetric binary systems with relatively large mass ratios. Specifically, we employ the open-source software PyCBC [53] and Bayesian inference methods [54, 55, 56] within the parameterized post-Einsteinian (ppE) framework [57, 58, 59, 60] to analyze GW data. Using the IMRPhenomXHM waveform model, our analysis of GW230529 data yields a constraint on the scalar-tensor coupling parameter of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25$. By including higher-mode effects we obtain constraints of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$. The former represents a 13.3% and 30.1% improvement over the dominant-mode case. Meanwhile, using the observations of GW230529, GW200115, and GW190814, which is assumed to be an NSBH event, the constraints improve further, such that $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$. This is currently the strongest constraint utilizing GWs.
The structure of this paper is as follows: Section 2, we explain the construction of the waveform models used in this paper. Section 3 describes the Bayesian inference method used for data analysis. Section 4 presents the main results, and Section 5 concludes the paper with a summary and outlook. Throughout, we adopt the convention $G=c=1$.
2 Waveform
In this section, we introduce the ppE formalism and show how the waveforms in screened modified gravity and Brans–Dicke theories map into it.
2.1 Parameter post-Einsteinian framework
The coalescence of compact binary systems can be categorized into three sequential phases: inspiral, merger, and ringdown. In the inspiral phase, the two compact objects remain well separated, moving at relatively low velocities, which permits the application of the post-Newtonian (PN) approximation to accurately model waveforms for systems with small mass ratios [61]. In our analysis, we employ the parameterized ppE framework—a general parametric method introduced by Yunes and Pretorius to characterize systematic deviations of PN waveforms in various modified gravity (MG) theories from those predicted by GR [62]. When considering only the leading-order correction, the inspiral waveform within the ppE formalism consists of the standard GR contribution supplemented by additional terms encoding possible modifications [57]:
$\tilde{h}(f)=\tilde{h}_{\rm GR}(f)\,(1+\alpha\,u^{a})e^{i\beta u^{b}}\,,$ (2.1)
where $f$ is the frequency, and $(\alpha u^{a},\beta u^{b})$ are the amplitude and phase corrections from modified gravity. $b=2{\rm~{}PN}-5$ and $a=b+5$ are the ppE order parameters, and the $\rm~{}PN$ in this equation symbolizes the PN order, $u={(\pi\mathcal{M}f)}^{1/3}$ is the reduced frequency of the inspiral that is proportional to the relative velocity of the binary, $\mathcal{M}=\eta^{3/5}M$ is the chirp mass, $M=m_{1}+m_{2}$ is the total mass, $\eta=m_{1}m_{2}/(m_{1}+m_{2})^{2}$ is the symmetrical mass ratio, $m_{1}$ and $m_{2}$ are the major and minor masses, respectively.
One can map Eq.(2.1) to the basis of spin-weighted spherical harmonics $Y_{-s}^{\ell m}(\iota,\varphi)$ to obtain the harmonic decomposition of ppE waveform [63]:
$\tilde{h}_{\rm ppE}(f)$ $=\sum_{\ell m}\tilde{h}_{\ell m}^{\rm ppE}(f)\,,$
$\tilde{h}_{\ell m}^{\rm ppE}(f)$ $=\tilde{h}^{\rm GR}_{\ell m}(f)\,(1+\alpha_{\ell m}\,u^{a}_{\ell m})e^{i\beta_{\ell m}u^{b_{\ell m}}}\,.$ (2.2)
In this paper, the GR waveform will be produced using IMRPhenomXHM, and the higher harmonics modes $(\ell,|m|)=\{(2,2),(2,1),(3,3),(3,2),(4,4)\}$ are considered. Both $\{a_{\ell m},\alpha_{\ell m},b_{\ell m},\beta_{\ell m}\}$ can be different for different values of $\ell$ and $m\,$.
We focus on the phase correction because of the negligibility of the amplitude correction. Hence the ppE waveform reduces to
$\tilde{h}_{\ell m}^{\rm ppE}(f)=\tilde{h}_{\ell m}^{\rm{GR}}(f)e^{i\beta_{\ell m}u^{b_{\ell m}}}.$ (2.3)
The phase correction with higher harmonics in (2.3) can be derived as [63]:
$\beta_{\ell m}=\beta_{m},\qquad b_{\ell m}=b,$ (2.4)
with
$\beta_{m}=\left(\frac{2}{m}\right)^{b/3-1}\beta_{22},$ (2.5)
2.2 Waveform in scalar-tensor theory
In this work we consider a class of scalar-tensor theories, which can be described by the [64]
$S=$ $\frac{1}{16\pi}\int d^{4}x\sqrt{-\tilde{g}}\left[\tilde{R}-2\tilde{g}^{\mu\nu}\left(\partial_{\mu}\varphi\right)\left(\partial_{\nu}\varphi\right)-V(\varphi)\right]$ (2.6)
$+S_{M}\left[\Psi,\phi(\varphi)\tilde{g}_{\mu\nu}\right]\,.$
Here, $\tilde{g}^{\mu\nu}$ and $\tilde{g}$ denote the Einstein-frame metric and its determinant, respectively. $\tilde{R}\equiv\tilde{g}^{\mu\nu}\tilde{R}^{\mu\nu}$ is the Ricci scalar. $\varphi$ denotes the scalar field, and $V(\varphi)$ is its potential, which gives the field an effective mass. For Brans–Dicke theory, the potential term $V(\varphi)$ is not considered. $\Psi$ denotes the matter fields and $\phi(\varphi)$ is the conformal coupling function. In scalar-tensor theories, compact binary systems exhibit a misalignment between the centers of gravitational binding energy and inertial mass. This asymmetry leads to a time-varying dipolar structure, which radiates additional energy through scalar emission [65]. In our analysis, we adopt a waveform model incorporating the leading-order correction to general relativity, characterized by a dipole contribution to the gravitational-wave phase [65, 60, 66, 67],
$h(f)=h_{\rm GR}(f)\exp\left[i\frac{3}{128\eta}\varphi_{-2}(\pi Mf)^{-7/3}\right].$ (2.7)
Here $\eta=m_{1}m_{2}/M^{2}$ is the symmetric mass ratio. $\varphi_{-2}=-5(\alpha_{\rm A}-\alpha_{\rm B})^{2}/168$, where $\alpha_{\rm A}\text{ and }\alpha_{\rm B}$ are scalar charges of the components of the binary. For black holes, the no-hair theorem prevents them to acquire scalar charges. In many scalar-tensor theories including the models considered in this work, where the no-hair theorem can be applied, scalar charges of black holes are 0 [68, 69, 70].
For screened modified gravity theory, we only consider the leading order modification and replace $\alpha_{\rm A}-\alpha_{\rm B}$ as $\varphi_{\rm{}_{VEV}}/({M_{\rm Pl}\Phi_{A}})$. Where $M_{\rm Pl}=\sqrt{1/8\pi G}$ is the reduced Planck mass, $\varphi_{\rm VEV}$ is the vacuum expectation value of the scalar field, and $\Phi_{A}=Gm/R$ is the surface gravitational potential of the object $A$. In conjunction with Eq.(2.1), the dominant-mode correction of these ppE parameters in screened modified gravity theory can be written as [43, 60, 67]:
$\beta_{22}^{\rm SMG}$ $=$ $\frac{-5}{7168}\eta^{2/5}\varphi_{\rm VEV}^{2}/({M_{\rm Pl}\Phi_{A}})^{2}.$ (2.8)
For Brans-Dicke theory, where one can replace $(\alpha_{\rm A}-\alpha_{\rm B})^{2}$ as $2(s_{\rm A}-s_{\rm B})^{2}/(2+\omega_{\rm BD})$ [43]. Here, $s_{\rm A}$ denotes the sensitivity of the $A$th body, which approximately equals its compactness (about 0.5 for black holes and $\sim 0.2$ for neutron stars). Combining the insights from Eq.(2.1), the dominant-mode correction of these ppE parameters in BD theory can be written as [43, 60, 67]:
$\beta_{22}^{\rm BD}$ $=$ $\frac{-5}{3584}\eta^{2/5}(s_{\rm A}-s_{\rm B})^{2}/(2+\omega_{\rm BD}).$ (2.10)
The leading order modification in the GW waveform introduced by scalar-tensor theory starts at the PN order, corresponding to $b^{\rm SMG}=b^{\rm BD}=-7$. And by combining Eqs.(2.4) and Eq.(2.5), we can derive $\beta_{\ell m}^{\rm SMG}$ and $\beta_{\ell m}^{\rm BD}$.
3 Bayesian inference
The ppE waveform formalism introduces a parametric model to describe possible deviations from general relativity in gravitational wave signals. This model can be incorporated into Bayesian inference frameworks to extract physical parameters or place constraints on modified gravity theories from observed waveforms [54, 55, 56]. Bayesian inference is based on Bayes’ theorem, which relates the conditional and marginal probabilities of random events. Given a model hypothesis $\mathcal{H}$ and a set of observational data $d$, the posterior probability distribution of the model parameters $\bm{\vartheta}$ can be obtained via Bayes’ theorem:
$p(\bm{\vartheta}|d,\mathcal{H})=\frac{p(\bm{\vartheta}|\mathcal{H})p(d|\bm{\vartheta},\mathcal{H})}{p(d|\mathcal{H})},$ (3.1)
where $p(\bm{\vartheta}|\mathcal{H})$ represents the prior distribution over parameters under the hypothesis $\mathcal{H}$, and $p(d|\bm{\vartheta},\mathcal{H})$ is the likelihood function, which quantifies the probability of observing the data $d$ assuming the parameters $\bm{\vartheta}$. The likelihood is not derived directly from the waveform model itself, but rather constructed by statistically comparing the observed data with theoretical waveform predictions—such as those generated under the parameterized post-Einsteinian (ppE) formalism—while accounting for the noise properties of the detectors. Typically, the noise is modeled as stationary and Gaussian. The denominator $p(d|\mathcal{H})$ is known as the Bayesian evidence, which serves as a normalization factor and encapsulates the overall plausibility of the hypothesis $\mathcal{H}$ given the data.
$m_{1}$ Mass of the major component
$m_{2}$ Mass of the minor component
$\chi_{1z}$ Aligned or anti-aligned spins of the major component
$\chi_{2z}$ Aligned or anti-aligned spins of the minor component
$\alpha_{S}$ Right ascension of source location
$\delta$ Declination of the source location
$\psi$ Polarization angle
$\iota$ Inclination angle
$\phi_{\rm ref}$ Phase at the reference frequency
$t_{c}$ Coalescence time
$D_{L}$ Luminosity distance
$\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}},\omega_{\rm{}_{BD}}$ Parameters to be bounded of SMG and BD separately.
Table 1: Summary of physical parameters and their meaning.
Under the assumption of stationary Gaussian noise, the logarithm of the likelihood function takes the following form:
$\log p(d|\bm{\vartheta},\mathcal{H})=\log\bar{\alpha}-\frac{1}{2}\sum_{k}\left<d_{k}-h_{k}(\bm{\vartheta})\middle|d_{k}-h_{k}(\bm{\vartheta})\right>,$ (3.2)
where $\log\bar{\alpha}$ is a normalization constant, the index $k$ labels individual detectors in the interferometer network, and $d_{k}$ and $h_{k}(\bm{\vartheta})$ denote the observed data and template waveform, respectively, for the $k$-th detector. The inner product is defined in the frequency domain as
[…]
where $S_{n}(f)$ denotes the one-sided power spectral density (PSD) of the detector noise, and $(\tilde{a}(f),\tilde{b}(f),\tilde{a}^{*}(f),\tilde{b}^{*}(f))$ are the Fourier transforms of the time-domain signals $a(t)$ and $b(t)$, respectively. $f_{\rm low}$ denotes the low-frequency cutoff of the gravitational wave data, $f_{\rm max}=f_{\rm ISCO}=(6^{3/2}\pi M)^{-1}$ represents the upper cutoff frequency of inspiral stage.
In our parameter estimation analysis, we sample over a total of 12 intrinsic and extrinsic parameters, including the parameter $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}$ in the screened modified gravity theory and Brans–Dicke coupling constant $\omega_{\rm BD}$. The physical interpretations of each parameter are summarized in Table 1.
4 Results
In scalar–tensor theories, corrections to the GW phase relative to GR depend on the difference in sensitivities of the compact binary components. Consequently, for symmetric binaries with equal sensitivities, the scalar–tensor correction vanishes, and the waveform is indistinguishable from GR predictions [43].
Recently, the LIGO-Virgo-KAGRA (LVK) collaboration reported the detection of GW230529, a GW signal most plausibly originating from the coalescence of a NSBH. Leveraging this event, we constrain scalar–tensor theories by performing Bayesian inference over the parameter space $\bm{\vartheta}$.
Figure 1: Probability density of Bayesian posterior distributions on $\varphi_{-2}$. The gray line corresponds to the constraint at a 90% credible level by using the dominant mode (DM) correction, while the red line stands for the constraint at the 90% confidence level by using the correction including the higher mode (HM).
The analysis is carried out using the open-source PyCBC software package [53], employing MCMC sampling with the emcee_pt sampler [71], initialized with 200 walkers. To mitigate contamination from low-frequency noise, we adopt a lower frequency cutoff of $f_{\rm low}=20$ Hz. A data segment of 64 seconds centered around GW230529 is used for parameter estimation. As for prior choices, we assume a uniform distribution for $1/(\omega_{\rm BD}+2)$ in the range $[0,0.5]$, and for $\varphi_{-2}$ in the interval $[-0.0027,0]$. Spin components $\chi_{1z}$ and $\chi_{2z}$ are aligned or anti-aligned with the orbital angular momentum, and are independently drawn from a uniform distribution in the range $[-0.99,0.99]$.
Figure 2: Probability density of Bayesian posterior distributions on $\omega_{\rm{}_{BD}}$. The gray line corresponds to the constraint on BD at 90% credible level by using the dominant mode (DM) correction, while the red line represents the constraint on BD with 90% probability by using the correction including higher mode (HM). We use a logarithmic scale due to the posterior distribution of $\omega_{\rm BD}$ spanning several orders of magnitude.
Our analysis focuses on the inspiral regime and neglects tidal interactions in NSBH systems [72], and we employ the IMRPhenomXHM waveform model from the LALSimulation package [73] for both the screened modified gravity and Brans–Dicke hypotheses. These phenomenological models operate in the frequency domain and incorporate the effects of non-precession and higher-order multipole moments.
Figure 3: Probability density of Bayesian posterior distributions on $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}$. The gray line corresponds to the constraint at a 90% credible level by using the dominant-mode (DM) correction, while the red line stands for the constraint at the 90% confidence level by using the correction including the higher mode (HM).
We present the resulting constraints on $\varphi_{-2}$ in scalar–tensor gravity in Fig. 1. For GW230529, we obtain a 90% credible lower bound of $\varphi_{-2}>-1.98\times 10^{-4}$ using the dominant mode (DM) correction only. When higher-order modes HM are included, the constraint improves to $\varphi_{-2}>-1.54\times 10^{-4}$, corresponding to a 22.22% enhancement. Corresponding constraints on the Brans–Dicke coupling parameter $\omega_{\rm BD}$ are shown in Fig. 2. For GW230529, the dominant-mode analysis yields a 90% lower bound of $\omega_{\rm BD}>25$. Including higher-order modes improves the bound to $\omega_{\rm BD}>32.68$, a 30.1% enhancement. For sreened modified gravity theory, we adopt neutron star parameters $m=1.4M_{\odot}$ and $R=10$km. Using the same Bayesian pipeline and the GW230529 signal, we constrain the scalar field vacuum expectation value $\varphi_{\rm VEV}$. We find that $\varphi_{\rm VEV}/M_{\rm Pl}<1.7\times 10^{-2}$ using the dominant mode, and $\varphi_{\rm VEV}/M_{\rm Pl}<1.5\times 10^{-2}$ when higher-order modes are included.
We also consider a combined analysis of the three NSBH events, GW230529, GW200115 and GW190814. Due to convergence issues in the dominant-mode-only case [51], we report only the result using higher-order modes for the combined dataset: $\varphi_{-2}>-4.26\times 10^{-5}$, demonstrating substantial improvement. In the combined analysis of the three NSBH events incorporating higher modes, the constraint strengthens significantly to $\omega_{\rm BD}>123.75$. Similarly, Considering the combination, we obtain a tighter constraint as well $\varphi_{\rm VEV}/M_{\rm Pl}<7.84\times 10^{-3}$. This represents the most stringent bound to date using gravitational-wave observations, assuming GW190814 is indeed an NSBH event. All the above resulting bounds are summarized in Table 2.
GW230529 Combination (GW230529,GW200115,GW190814)
Dominant mode (DM) Higher mode (HM) Higher mode (HM)
$\varphi_{-2}$ $-1.98\times 10^{-4}$ $-1.54\times 10^{-4}$ $-4.26\times 10^{-5}$
$\frac{\varphi_{\rm VEV}}{M_{\rm Pl}}$ $1.7\times 10^{-2}$ $1.5\times 10^{-2}$ $7.84\times 10^{-3}$
$\omega_{\rm BD}$ 25.12 32.68 123.75
Table 2: Summary of constraints on scalar–tensor gravity at a 90% credible level.
5 Conclusion and discussion
In this work, we revisited gravitational wave generation in scalar–tensor theories and outlined the construction of waveform models within the parameterized post-Einsteinian (ppE) framework. Using Bayesian inference with MCMC sampling implemented via the PyCBC software package, we analyzed the event GW230529 and derived constraints on screened modified gravity and Brans–Dicke theory. Motivated by the importance of higher-order multipole modes in asymmetric binaries with large mass ratios, we employed the IMRPhenomXHM waveform model to incorporate such effects. Our analysis yields bounds of $\varphi_{\rm VEV}/M_{\rm Pl}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25$ using only the dominant mode, which improve to $\varphi_{\rm VEV}/M_{\rm Pl}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$ when higher modes are included, representing 13.3% and 30.1% enhancements, respectively. A joint analysis of GW230529, GW200115, and GW190814 (assumed to be a neutron star–black hole event) further tightens the constraints to $\varphi_{\rm VEV}/M_{\rm Pl}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$, currently the most stringent limits obtained from gravitational wave observations.
It should be noted that our treatment of the PSD differs from that of others, who use the available event-specific PSDs provided in the LVK posterior sample releases, whereas we estimate the PSD directly from the strain data using the Welch method, which more closely resembles a real event [53]. Second, the choice of the upper cutoff frequency $f_{\rm max}=f_{\rm ISCO}=(6^{3/2}\pi M)^{-1}$ is not treated as a fixed value in our analysis but rather varies across different MCMC realizations. As a result, the constraints we obtain may be more conservative compared to others but are also more rigorous and faithful to the actual data.
Given that meaningful constraints on scalar-tensor gravity require observations of NSBH binaries, and the current number of confirmed NSBH events remains limited, the scope for testing such theories is still narrow. However, the prospects are promising: future gravitational wave observatories, both ground-based and space-based, including next generation detectors are expected to significantly increase the number of detected NSBH events. This will enable more stringent and statistically robust constraints on scalar-tensor modifications of gravity. Furthermore, we anticipate the detection of systems with higher mass ratios, for which higher-order effects become increasingly prominent. These systems will enhance the sensitivity of gravitational wave measurements to deviations from general relativity, enabling more precise modeling of waveforms and stronger tests of alternative theories.
Acknowledgments
This work is supported by “the Fundamental Research Funds for the Central Universities” under the reference No. 2042024FG0009. The data, software, and/or web tools used in this research were obtained from the Gravitational Wave Open Science Center, a service of LIGO Laboratory, the LIGO Scientific Collaboration, and the Virgo Collaboration.